Hubert and Kate are farmers, each one owns a 12 acre plot of land. The following table shows the amount of corn and rye each farmer can produce per year on a given acre. Each farmer choose whether to devote all acres to producing corn or rye or to produce corn on some of the land and rye on the rest.

Corn (Bushels per acre) Rye (Bushels per acre)

Hubert 12 3

Kate 15 5

(a) Plot the PPF curve for Hubert and kate

(b) What is Hubert’s opportunity cost of producing 1 bushel of rye?

(c) What is the Kate’s opportunity cost of producing 1 bushel of rye?


1
Expert's answer
2021-05-30T14:18:30-0400


The vertical axis measures the quantity of corn in bushels per acre and the horizontal axis represents the quantity of rye in bushels per acre.

Hubert could either produce 12 bushels of corn or 3 bushels of rye per acre or combinations of two goods produced per acre that are traced by the Hubert's PPF line.

Kate could either produce 15 bushels of corn or 5 bushels of rye per acre or combinations of two goods produced per acre that are traced by the Kate's PPF line. 


(b)

Hubert's opportunity cost in producing a unit of rye


"=\frac{Bushels\space of\space corn\space production\space forgone}{Bushels\space of\space rye\space produced}"


"=\frac{12}{3}=4"

Hubert’s opportunity cost of producing 1 bushel of rye is equal to 4.


(c)


Kate's opportunity cost in producing a unit of rye


"=\frac{Bushels\space of\space corn\space production\space forgone}{Bushels\space of\space rye\space produced}"


"=\frac{15}{5}=3"


Kate’s opportunity cost of producing 1 bushel of rye is equal to 3.